# FlowSearch: Advancing deep research with dynamic structured knowledge flow

### v1

@@ Abstract
Deep research is an inherently challenging task that demands both breadth and depth of thinking. It involves navigating diverse knowledge spaces and reasoning over complex, multi-step dependencies, which presents substantial challenges for agentic systems. To address this, we propose FlowSearch, a multi-agent framework that actively constructs and evolves a dynamic structured knowledge flow to drive subtask execution and reasoning. FlowSearch is capable of strategically planning and expanding the knowledge flow to enable parallel exploration and hierarchical task decomposition, while also adjusting the knowledge flow in real time based on feedback from intermediate reasoning outcomes and insights. FlowSearch achieves state-of-the-art performance on both general and scientific benchmarks, including GAIA, HLE, GPQA and TRQA, demonstrating its effectiveness in multi-disciplinary research scenarios and its potential to advance scientific discovery. The code is available at https://github.com/Alpha-Innovator/InternAgent.
Figure 1: FlowSearch (Ours) achieves leading performance on the GAIA, GPQA, and HLE benchmarks, outperforming competitive agent workflow methods (OpenAI-DeepResearch, MiroFlow, Manus, OWL) as well as LLM-based approaches (GPT-5, Intern-S1, DeepSeek-R1).
@@ 1 Introduction
The rapid progress of Large Language Models (LLMs) [1, 2, 3, 4] has marked a significant milestone in artificial intelligence, exhibiting impressive capabilities in natural language understanding, generation, and complex reasoning [5, 6]. Researchers have found that the capabilities of large language models (LLMs) can be harnessed to build agent systems for handling diverse tasks [7, 8]. More notably, LLMs are proving to be valuable tools in facilitating scientific research and discovery [9, 10]. However, effectively leveraging these capabilities in open-ended research contexts remains a non-trivial challenge—demanding not only iterative hypothesis formulation and strategic information acquisition, but also the orchestration of multi-step reasoning within dynamic and often uncertain knowledge spaces. These considerations have motivated the development of Deep Research (DR) systems—frameworks that transcend isolated reasoning or passive retrieval by integrating LLMs within structured, goal-directed workflows. The design of such systems is essential for unlocking the full potential of LLMs in enabling systematic scientific discovery across various domains.
Existing deep research systems [11, 12] often draw inspiration from either individual or collaborative research paradigms. (1) Single-agent paradigm: [13, 14, 15, 5] LLMs centrally manage the research workflow by leveraging a long context window to accumulate and reason over information. While this setup mirrors the behavior of an individual researcher, it is prone to tunnel vision—overcommitting to early hypotheses and lacking the breadth necessary for wide-ranging exploration. (2) Multi-agent paradigm: [11, 16, 9] such designs scale up research efforts by leveraging explicit planning and role specialization among agents. However, the prevalent reliance on serial plan execution in these systems places strict demands on context management. Excessively retrieved information may result in context window overflow and easily overwhelm the system [17], thereby reducing the system’s ability to sustain deep and coherent reasoning. Overall, both paradigms face inherent trade-offs between exploratory breadth and reasoning depth, highlighting the need for more adaptive, reflective, and context-sensitive research agents.
In this work, we present FlowSearch, a multi-agent system built upon a dynamic knowledge flow framework that enables structured and efficient knowledge propagation throughout the process of scientific discovery. The system begins with a flow planner that constructs an initial knowledge flow, where nodes represent subproblems to be solved or key concepts to be retrieved, and edges encode the knowledge dependencies among them. As discovery progresses, the knowledge flow can be incrementally expanded to ensure both exploratory breadth and reasoning depth, while its structure remains dynamically revisable in response to intermediate findings—enabling the overall process to proceed in a reflective and adaptive manner. Importantly, each node in the flow not only guides the execution of subtasks but also supports recursive decomposition, integration of upstream knowledge, and local summarization of intermediate results. This leads to a more refined and contextually relevant knowledge stream, which allows for deep reasoning within local regions of the flow while maintaining global coherence through dynamic flow-level adjustments. Such a design ensures the efficiency of the knowledge flow and enhances the system’s ability to perform complex, multi-step scientific problem solving.
@@
To demonstrate the superior performance of FlowSearch, we conduct experiments on several challenging benchmarks, including GAIA [18], which evaluates the general problem-solving abilities of AI assistants, as well as three scientific-question-answering benchmarks HLE [19],GPQA [20] and TRQA [10]. FlowSearch achieves state-of-the-art performance on GAIA, HLE and TRQA, and demonstrates highly competitive results on GPQA. These findings highlight the strong problem-solving capability of FlowSearch, enabled by the integration of graph-driven planning. In summary, our main contribution can be described as follows:
  • We propose a novel dynamic structured knowledge flow to encode the logic in complex problem solving, enabling the deep research agent to explicitly capture dependencies among subproblems and key concepts, in contrast to conventional frameworks.
@@
  • We develop FlowSearch, a novel multi-agent system for deep research built upon the dynamic structured knowledge flow, capable of generating structured plans and dynamically refining them during execution to enhance performance.
@@
  • We evaluate FlowSearch on the general AI assistant benchmark GAIA and the multi-disciplinary scientific question-answering benchmarks HLE, GPQA and TRQA, demonstrating state-of-the-art performance across all of them.
@@ 2 FlowSearch
FlowSearch is built upon a dynamic knowledge flow that enables structured and adaptive scientific research. As illustrated in Figure 2, the system comprises three core components: Knowledge Flow Planner, which constructs high-quality knowledge flows tailored to the research objective; Knowledge Collector, which executes subtasks and enriches each node with relevant contextual information; and Knowledge Flow Refiner, which monitors progress and dynamically adjusts the flow based on intermediate outcomes and newly acquired knowledge. By enabling multiple agents to collaborate along this evolving flow, FlowSearch achieves systematic, scalable, and efficient problem solving. We begin by formalizing the concept of the structured knowledge flow, followed by detailed descriptions of Knowledge Flow Planner, Knowledge Collector, and Knowledge Flow Refiner.
@@ 2.1 Structured Knowledge Flow
Structured Knowledge Flow provides principled guidance for systematically organizing information to improve both the systematicity and effectiveness of deep research.
A common workflow in deep research agents is to address a user query $q$ by assembling a strictly linear pipeline $L(q)=[s_{1},s_{2},\dots,s_{n}]$ and executing it in order $s_{1}\!\to\!s_{2}\!\to\!\cdots\!\to\!s_{n}$. The precedence relations are implicitly encoded by positional order, i.e., $s_{i}\prec s_{j}\iff i<j$. Despite its procedural simplicity and ease of implementation, such a linear formalism fails to capture the inherently complex and non-linear dependencies of real-world research processes.
@@
By formalizing the research process in this manner, we enable the agent not only to generate execution plans but also to reason over the structural dependencies among subtasks, ensuring coherence and systematicity throughout the deep research workflow.
@@ 2.3 Knowledge Collector
where $t_{i}$ and $d_{i}$ are the task type and task description of node $v_{i}$, $f^{exec}(\cdot)$ is the LLM executor with tools depending on the context knowledge $\{c_{j}|(v_{j}\rightarrow v_{i})\in E_{t}\}$. After the parallel execution of all the outermost executable nodes, a flow refinement will be conducted based on the newly obtained knowledge, which will be detailed as follows.
@@ 2.5 Conclusion Generation
At the conclusion of all iterations of knowledge collector and flow refiner, only the initial query node remains unexecuted. If the query is a scientific question that can be answered directly and simply, the query node will be executed using the knowledge from the connected nodes. If the query is to create a detailed scientific report, the final query node will gather knowledge from all nodes in the flow, complete the reasoning process, and provide a full report.
@@ 3.1 Experiments Setup
  • GAIA [18]: a benchmark of real-world questions that require a set of fundamental abilities such as reasoning, multi-modality handling, web browsing, and generally tool-use proficiency. Our results are based on its 165-question validation set.
  • GPQA [20]: a benchmark of 448 multiple-choice questions across biology, chemistry, and physics, authored by domain experts to ensure depth and rigor, thereby providing a stringent evaluation of advanced reasoning and scientific knowledge. We use its 198-question GPQA-diamond subset for evaluation.
  • HLE [19]: Humanity’s Last Exam is a multimodal benchmark consisting of 2,500 questions across mathematics, humanities, and natural sciences. Developed by subject experts, it provides a frontier-level test of academic competence where current LLMs still perform far below human experts.
  • TRQA [10]: a domain-specific benchmark for therapeutic target discovery. It covers fundamental biology, disease biology, pharmacology, and clinical medicine, providing a systematic evaluation framework for biomedical research agents. We use its 172-question TRQA-lit subset for evaluation.
Methods of Comparison. To validate the effectiveness of FlowSearch, we compare FlowSearch on GAIA, GPQA, HLE and TRQA against both cutting-edge large language models including Qwen3 series model [3], Intern-S1 [4], Deepseek-R1 [1], GPT-o4-mini and GPT-5 [2], and some state-of-the-art deep research agent, including open-source frameworks OpenAI-DR (Deep Research) [21], Gemini-DR [22] and Manus [16], and proprietary approaches MiroFlow [12], OWL [11] X-Masters [23] and Origene [10]. In the experiments, we prompt GPT-o4-mini to serve as the Knowledge Flow Planner, Knowledge Collector and Knowledge Flow Refiner in our workflow.
Table 1: Performance comparison on GAIA, GPQA-diamond and HLE benchmarks. The best results are bolded and the second best results are underlined. Results not reported in the original papers are denoted as “ - ".
@@
Method Level 1 Level 2 Level 3 Avg. Bio Chem Phys Avg.
@@
Qwen-3-8B 11.32 2.32 0.00 4.85 - - - 44.44 - -
Qwen3-32B 13.21 3.49 3.84 6.67 - - - 49.49 - -
Qwen3-235B 15.09 3.49 3.84 7.27 - - - 47.47 9.18 8.60
Intern-S1 28.30 9.30 7.69 15.15 89.47 59.49 93.02 78.26 8.90 8.30
Deepseek-R1 33.96 13.95 3.84 18.78 63.16 76.34 91.86 82.32 8.60 -
o4-mini 28.30 12.79 7.69 16.97 78.95 63.44 94.19 78.28 14.50 14.28
GPT-5 - - - - 84.21 76.34 95.35 85.35 25.85 24.76
@@
OpenAI DR 74.29 69.06 47.60 67.36 - - - - - 26.60
Manus 86.50 70.10 57.70 73.30 - - - - - -
Gemini Deep Research - - - - - - - - - 26.90
@@
MiroFlow - - - 74.50 - - - - 29.50 27.20
OWL 84.90 68.60 42.30 69.70 - - - - - -
X-Masters - - - - - - - - 32.10 27.72
@@
FlowSearch (Qwen-235B) 69.81 60.47 30.77 58.79 63.16 58.06 75.58 66.16 15.04 14.84
FlowSearch (o4-mini) 90.56 76.74 50.00 76.96 84.21 79.57 96.51 87.37 31.60 30.80

### v2

@@ Abstract
Deep research is an inherently challenging task that demands both breadth and depth of thinking. It involves navigating diverse knowledge spaces and reasoning over complex, multi-step dependencies, which presents substantial challenges for agentic systems. To address this, we propose FlowSearch, a multi-agent framework that actively constructs and evolves a dynamic structured knowledge flow to drive subtask execution and reasoning. FlowSearch is capable of strategically planning and expanding the knowledge flow to enable parallel exploration and hierarchical task decomposition, while also adjusting the knowledge flow in real time based on feedback from intermediate reasoning outcomes and insights. FlowSearch achieves competitive performance on both general and scientific benchmarks, including GAIA, HLE, GPQA and TRQA, demonstrating its effectiveness in multi-disciplinary research scenarios and its potential to advance scientific discovery. The code is available at https://github.com/InternScience/InternAgent.
@@ 1 Introduction
The general capabilities of Large Language Models (LLMs) enable agent systems for diverse tasks, supporting scientific research and discovery [undeff, undefg, undefh, undefi]. However, effectively utilizing these capabilities in open-ended research requires iterative hypothesis formulation, strategic information acquisition, and multi-step reasoning in dynamic and uncertain knowledge spaces. These challenges have inspired the development of Deep Research (DR), which combines LLMs with expert-designed knowledge to go beyond basic reasoning or simple information retrieval. Such systems are essential for unlocking the potential of LLMs in enabling scientific discovery on various domains.
Existing deep research systems [undefj, undefk] are primarily inspired by either individual or collaborative research paradigms. (1) Single-agent paradigm: [undefl, undefm, undefn, undefd] a single LLM centrally manages the research workflow using a long context window to accumulate and reason over information. While this mirrors individual researchers, it is prone to tunnel vision, overcommitting to early hypotheses and limiting exploratory breadth. (2) Multi-agent paradigm: [undefj, undefo, undefh] research is scaled via explicit planning and role specialization. However, serial plan execution requires maintaining context across multiple planning steps and agents, forcing aggressive context truncation or retrieval filtering. As a result, intermediate reasoning chains are often fragmented, leading to shallow, step-wise reasoning and a loss of global reasoning depth [undefp]. Overall, both paradigms trade off exploratory breadth against reasoning depth, motivating the need for more adaptive and context-aware research agents.
In this work, we propose a novel Dynamic Structured Knowledge Flow to enable structured and efficient knowledge propagation throughout scientific discovery activities, and instantiate it in FlowSearch, a multi-agent system grounded in this design. Unlike conventional sequential planning, the proposed Structured Knowledge Flow models scientific discovery as an evolving graph (as illustrated in Fig. 4) of interdependent knowledge units, where exploratory breadth and reasoning depth are jointly supported through incremental expansion and structural revision. Within this flow, knowledge is decomposed, integrated, and summarized at different granularity levels, enabling deep local reasoning while preserving global coherence.
FlowSearch instantiates this abstraction as a multi-agent system by operationalizing the dynamic structured knowledge flow throughout the research process. It starts with a flow planner that initializes a graph-structured knowledge flow, where nodes represent subproblems or key concepts and edges capture their knowledge dependencies. As scientific discovery progresses, the knowledge flow is continuously expanded and dynamically revised in response to intermediate findings. Each node orchestrates the execution of corresponding subtasks while supporting recursive decomposition, integration of upstream knowledge, and local summarization, thereby enabling structured and efficient knowledge propagation with effective context management across complex, multi-step scientific problem solving.
Figure 1: FlowSearch (Ours) achieves leading performance on the GAIA, GPQA, and HLE benchmarks, outperforming competitive agentic frameworks (OpenAI-DR, MiroFlow, Tongyi-DR and Manus) as well as LLM-based approaches (GPT-5, DeepSeek-R1, MiroThinker).
@@
We conduct experiments on several challenging benchmarks, including GAIA [undefq], which evaluates the general problem-solving abilities of AI assistants, as well as three scientific-question-answering benchmarks HLE [undefr], GPQA [undefs] and TRQA [undefi]. As shown in Fig. 1, our approach achieves state-of-the-art results on GAIA, HLE, TRQA and GPQA. These findings highlight the strong problem-solving capability of FlowSearch, enabled by the integration of graph-driven planning. In summary, our main contribution can be described as follows:
  • Unlike conventional sequential frameworks in deep research agents, we introduce a novel dynamic structured knowledge flow to achieve a trade-off between exploratory breadth and reasoning depth.
@@
  • We develop FlowSearch, a multi-agent system built upon the dynamic structured knowledge flow, capable of generating structured plans and dynamically refining them during execution phase to enhance performance.
@@
  • We evaluate FlowSearch on the general AI assistant benchmark GAIA and the multi-disciplinary scientific benchmarks HLE, GPQA and TRQA, demonstrating state-of-the-art results.
2 Related Work
2.1 Agentic Systems
Agentic systems with LLM have evolved from static prompting to perception–action loops, enabling systems to plan [undefw], act [undefd], and learn using external tools. Foundational approaches, such as interleaved reasoning–acting frameworks [undefd] and tree search planning [undefx], improve reliability in multi-step tasks, while reflective self-revision mechanisms [undefy] and external memory [undefz] enhance long-horizon consistency. Recent efforts like OpenHands [undefaa] further expands action spaces and mitigate hallucinations, and evaluates on more realistic interactive benchmarks including AgentBoard [undefab], StuLife [undefac], and SWE-bench Verified [undefad]. Multi-agent orchestration involves role-specialized collaboration and negotiation [undefae, undefaf], replacing single-agent end-to-end optimization with a modular and scalable approach.
Despite these advances, most general-purpose agents target short to medium horizon tasks and interactive environments. Scientific research [undeft], however, requires handling long-horizon workflows, integrating diverse knowledge, and adapting strategies dynamically. This motivates the development of research-oriented agents, which focus on structured, adaptive, and knowledge-driven scientific inquiry.
2.2 Deep Research Agents
Recent advances in DR agents extend LLMs from retrieval-augmented generation to dynamic, tool-driven research workflows. Early systems such as WebGPT [undefag] and Toolformer [undeff] explored web and API integration, while industrial solutions e.g., OpenAI DR [undeft], Gemini DR [undefu], Grok DR [undefah], Perplexity DR [undefai], incorporate adaptive planning, iterative retrieval, and multimodal reasoning. Recently, single-agent designs (e.g., Search-o1 [undefal], WebDancer [undefl], Tongyi DeepResearcher [undefam]) enable end-to-end optimization, while multi-agent architectures (e.g., AI Scientist [undefaj], Agent Laboratory [undefak], and InternAgent [undefh]) offer modularity and scalability—crucial for complex research.
Recent studies, e.g., GeAR [undefap], PANGU DeepDiver [undefaq], Alita [undefaf], also show the benefit of explicit structures and self-evolving mechanisms for multi-hop reasoning. However, the existing DR agents still suffer from sequential bottlenecks and limited hierarchical decomposition, motivating frameworks like our FlowSearch that integrate multi-agent coordination with dynamic structured knowledge flow.
@@ 3 FlowSearch
FlowSearch is built upon a dynamic knowledge flow that enables structured and adaptive scientific research. As illustrated in Fig. 2, the system comprises three core components: Knowledge Flow Planner, which constructs high-quality knowledge flows tailored to the research objective; Knowledge Collector, which executes subtasks and enriches each node with relevant contextual information; and Knowledge Flow Refiner, which monitors progress and dynamically adjusts the flow based on intermediate outcomes and newly acquired knowledge. By enabling multiple agents to collaborate along this evolving flow, FlowSearch achieves systematic, scalable, and efficient problem solving. We begin by formalizing the concept of the structured knowledge flow, followed by detailed descriptions of Knowledge Flow Planner, Knowledge Collector, and Knowledge Flow Refiner.
@@ 3.1 Structured Knowledge Flow
Structured Knowledge Flow provides principled guidance for systematically organizing information to improve both the systematicity and effectiveness of deep research. Moreover, its graph-structured formulation enables flexible revision of local research objectives and knowledge dependencies when necessary.
A common practice in deep research agents is to address a user query $q$ by assembling a strictly linear pipeline $L(q)=[a_{1},a_{2},\dots,a_{n}]$ and executing it in order $a_{1}\!\to\!a_{2}\!\to\!\cdots\!\to\!a_{n}$. The precedence relations are implicitly encoded by positional order, i.e., $a_{i}\prec a_{j}\iff i<j$. Despite its procedural simplicity and ease of implementation, such a linear formalism fails to capture the inherently complex and non-linear dependencies of real-world research processes.
@@
By formalizing the research process in this manner, the agent is enabled not only to generate execution plans, but also to reason over the structural dependencies among subtasks, ensuring coherence and systematicity throughout the deep research workflow. Moreover, the graph-structured formulation allows multiple non-dependent nodes to be executed in parallel, and enables flexible revision of local research components without interfering with other ongoing processes, including the insertion of new nodes or the modification of suboptimal planning decisions when errors are identified, thereby enhancing the robustness of the overall system. These mechanisms will be further elaborated in the subsequent section on the construction of FlowSearch.
@@ 3.3 Knowledge Collector
where $t_{i}$ and $d_{i}$ are the task type and task description of node $v_{i}$, $f^{exec}(\cdot)$ is the LLM executor with tools depending on the context knowledge $\{c_{j}|(v_{j}\rightarrow v_{i})\in E_{t}\}$. After the parallel execution of all the outermost executable nodes, a flow refinement will be conducted based on the newly obtained knowledge, which will be described in Sec. 3.4.
@@ 3.5 Conclusion Generation
At the end of all iterations of knowledge collector and flow refiner, only the initial query node remains unexecuted. If the query is a straightforward scientific question that can be answered directly and simply, the query node will utilize the knowledge from its connected nodes to provide an immediate response. If the query requires generating a detailed scientific report, the final query node will aggregate knowledge from all nodes within the flow, perform a comprehensive reasoning process, and deliver a complete and thorough report. The details are explained in Appendix D.2.
@@ 4.1 Experiments Setup
  • GAIA [undefq]: a benchmark of real-world questions that require a set of fundamental abilities such as reasoning, multi-modality handling, web browsing, and generally tool-use proficiency. Our results are based on its 165-question validation set.
  • GPQA [undefs]: a benchmark of 448 multiple-choice questions across biology, chemistry, and physics, authored by domain experts to ensure depth and rigor, thereby providing a stringent evaluation of advanced reasoning and scientific knowledge. We use its 198-question GPQA-diamond subset for evaluation.
  • HLE [undefr]: Humanity’s Last Exam is a multimodal benchmark consisting of 2,500 questions across mathematics, humanities, and natural sciences. Developed by subject experts, it provides a frontier-level test of academic competence where current LLMs still perform far below human experts.
  • TRQA [undefi]: a domain-specific benchmark for therapeutic target discovery. It covers fundamental biology, disease biology, pharmacology, and clinical medicine, providing a systematic evaluation framework for biomedical research agents. We use its 172-question TRQA-lit subset for evaluation.
Methods of Comparison. To validate the effectiveness of FlowSearch, we compare FlowSearch on GAIA, GPQA, HLE and TRQA against cutting-edge large language models including Qwen3 series model [undefb], Intern-S1 [undefc], Deepseek-R1 [undef], GPT-o4-mini, and GPT-5 [undefa], and some state-of-the-art deep research agent, including proprietary approaches OpenAI-DR (Deep Research) [undeft], Gemini-DR [undefu] and Manus [undefo], leading react agentic models MiroThinker [undefk], Tongyi-DR [team2025tongyi], WebShaper [undefm], WebDancer [undefl], and open-source frameworks MiroFlow [undefk], OWL [undefj], X-Masters [undefv], AWorld [undefao], and OriGene [undefi]. In the experiments, we utilize GPT-o4-mini to serve as the Knowledge Flow Planner, Knowledge Collector and Knowledge Flow Refiner in our workflow.
Table 1: Performance comparison on GAIA, GPQA-diamond and HLE benchmarks. The best results are bolded and the second best results are underlined. Results not reported in the original papers are denoted as “ - ". Note that some approaches, such as MiroThinker and TongyiDR, can only report results on the HLE text-only subset as these approaches lack multimodal ability.
@@
Method Base Model Level 1 Level 2 Level 3 Avg. Bio Chem Phys Avg.
@@
Qwen-3-8B - 11.32 2.32 0.00 4.85 - - - 44.44 - -
Qwen3-32B - 13.21 3.49 3.84 6.67 - - - 49.49 - -
Qwen3-235B - 15.09 3.49 3.84 7.27 - - - 47.47 9.18 8.60
Intern-S1 - 28.30 9.30 7.69 15.15 89.47 59.49 93.02 78.26 8.90 8.30
Deepseek-R1 - 33.96 13.95 3.84 18.78 63.16 76.34 91.86 82.32 8.60 -
o4-mini - 28.30 12.79 7.69 16.97 78.95 63.44 94.19 78.28 14.50 14.28
GPT-5 - - - - - 84.21 76.34 95.35 85.35 25.85 24.76
@@
OpenAI DR - 74.29 69.06 47.60 67.36 - - - - - 26.60
Manus - 86.50 70.10 57.70 73.30 - - - - - -
Gemini DR - - - - - - - - - - 26.90
React Agentic Model
WebDancer QwQ-32B 61.5 50.0 25.0 51.5 - - - - - -
WebShaper Qwen2.5-72B 69.2 63.4 16.6 60.1 47.37 52.69 81.40 64.65 - -
MiroThinker Qwen2.5-72B - - - 81.9 84.21 75.27 95.35 84.85 37.7 -
Tongyi DR Qwen3-30B-A3B - - - 70.9 78.95 67.74 95.35 80.30 32.9 -
@@
MiroFlow Claude-3.7 - - - 74.50 - - - - 29.50 27.20
OWL Gemini-2.5-Pro 84.90 68.60 42.30 69.70 57.89 61.29 86.05 71.72 - -
X-Masters Deepseek-R1 - - - - 78.95 68.82 94.19 80.81 32.10 27.72
JoyAgent Claude-4 86.79 77.91 42.31 75.15 78.95 65.59 91.86 77.27 - -
AWorld Gemini-2.5-Pro 88.68 77.91 53.85 77.58 73.68 66.67 93.02 78.79 - -
@@
FlowSearch Qwen3-235B 69.81 60.47 30.77 58.79 63.16 58.06 75.58 66.16 15.04 14.84
FlowSearch o4-mini 92.45 82.56 61.54 82.42 84.21 79.57 96.51 87.37 36.10 34.52